OpenStudy (anonymous):

Alex Stocks Up For Winter. He Buys 36 Cans Of Vegetables. He Pays 80 Cents Per Can For Tomatoes And 40 Cents Per Can For Corn, For A Total Cost Of 20.80 How Many Cans Of Tomatoes Does He Buy .

OpenStudy (anonymous):

@DamnImBad today i m too tired i will do it in a bit.

OpenStudy (anonymous):

what, was your text written in like 1964?

OpenStudy (anonymous):

put say \(x\) as the number of cans of tomatoes , then \(24-x\) is the number of cans of corn solve \[80x+40(24-x)=1280\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

you need so solve \[80x+40(24-x)=1280\] for \(x\) that will be the number of cans of tomatoes

OpenStudy (anonymous):

\[80x+960-40x=1280\] \[40x+960=1280\] \[40x=320\] one more step and you are done

OpenStudy (anonymous):

@DamnImBad it looks like satellite was using different numbers, but the concept is the same. Are you still confused?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Ok my turn to help. Lets pull apart the problem to come up with equations to help us solve

OpenStudy (anonymous):

There is a total amt of can, 36. That consists of c number of corn and t number of tomato cans correct?

OpenStudy (anonymous):

so our first equation is c + t = 36